OpenStudy (anonymous):

Simplify, step by step please: [(1/(x+h)^3) - (1/x^3)] / 15h

OpenStudy (anonymous):

this is horrid...

OpenStudy (anonymous):

Agreed.

OpenStudy (anonymous):

are you talking about my question domian and range

OpenStudy (anonymous):

you have not given a function there prodgers

OpenStudy (anonymous):

the domain andrange of f (-2) of f(1)

OpenStudy (anonymous):

for this one you need to make a common denominator for x^3 and (x+h)^3

OpenStudy (anonymous):

so what would be the answer I am so confused with this mess. I have not taken this kind of mathe in 20 yrs

OpenStudy (anonymous):

\[(x^3 - (x+h)^3 )/(x^3 * (x+h)^3)\]

OpenStudy (anonymous):

prodgers2, go back to your question and look at what I typed....

OpenStudy (anonymous):

so is that the range or the donmain

OpenStudy (anonymous):

by the way is this for a limit?

OpenStudy (anonymous):

Nope, no limit

OpenStudy (anonymous):

and the above is still divided by 15h (x+h)^3=x^3+3x^2h+3xh^3+h^3

OpenStudy (anonymous):

so the x^3 will cancel

OpenStudy (anonymous):

and the h too

OpenStudy (anonymous):

\[(3x^2+3xh+h^2)/(15(x^3 * (x+h)^3)\]

OpenStudy (anonymous):

oh no, there is a minus sign in front of this

OpenStudy (anonymous):

yes and it is asking to write the answer in interval notation

OpenStudy (anonymous):

and I dont see how you can simplify it further

OpenStudy (anonymous):

\[[(1/(x+h)^3) - (1/x^3)] / 15h = 0 \] ???

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

akiroso is that the range or domain?

OpenStudy (anonymous):

prodgers2, it's neither, this is not your problem.

OpenStudy (anonymous):

okkie dookie

OpenStudy (anonymous):

THANKS ANDRAS!!! XD